# A Survey of Robotic Language Grounding: Tradeoffs Between Symbols and Embeddings

### v1

@@ 2.4 Predefined Skills
Recent methods have used LLMs as a planner to map language commands to a sequence of predefined skills. Skills can be learned from data or manually specified. These predefined skills are a formal representation since they are discrete and manually specified. (Even if skills are learned, they are frequently learned from human-provided demonstrations, which provide the discrete structure for the skills.) They are also action-based representations specifying actions to take rather than goals or resultant states. For example, a robot might be provided with skills such as “pick up,” “put down,” “drive to refrigerator,” “drive to microwave,” and “clean table.” Then, the challenge is to map a natural language instruction such as “Get me the apple” and “Clean the table with the sponge” to a sequence of skills. Huang et al. (2022a) iteratively prompted an LLM to decompose a high-level task specification in natural language to a sequence of action descriptions. They used sentence similarities to map the proposed action descriptions to actions available in a simulated environment where their corresponding predefined low-level controllers can be executed. SayCan (Ichter et al., 2022) iteratively prompted an LLM to sequence pre-trained skills described by predefined verb phrases based on their probabilities of success from the current state to solve mobile manipulation tasks specified by natural language on a physical robot. Their approach relies on both formal representations and high-dimensional, end-to-end differentiable representations. Despite using a fixed ontology of skills, these are implemented using a deep approach. SayCan learns these skills using a multi-task value function conditioned on language. CAPE (Raman et al., 2024) also used an LLM planner to sequence predefined skills. When a plan does not meet the precondition of some skill, it re-prompted the LLM with corrective feedback. In addition to using an LLM to ground language to predefined skills, Inner Monologue (Huang et al., 2022b) also enabled language feedback based on pre-trained perception models for describing the scene and detecting successful execution of skills. In these methods, the specific skills are often given relatively little attention, yet having the right set of skills at the right level of granularity is critical to success. We categorize this method on the formal side of the spectrum because the skills impose significant structure. That said, it imposes less structure than goal-based methods such as logical representations, leaving the LLM to keep track of higher-level constructs such as constraints, sequencing, and conditionals.
@@ 3 Mapping from Natural Language to High-Dimensional Vectors to Actions
On the other end of the spectrum are approaches that map natural language instructions to high-dimensional, non-formal representations. These deep approaches are largely defined by data and learning, whereas formal, symbolic representations are largely human-crafted and invoke manually provided structure. The large models that enable deep and many symbolic approaches are created by self-supervised pretraining on large datasets scraped from the web. These models thereby learn a generative representation of images, text, and other modalities that capture background and task knowledge that is useful for robotic applications, for example, as in  (Driess et al., 2023). By learning to model a large and varied distribution, the models also learn to perform tasks useful for robotic language grounding, including translation, semantic parsing, and broadly useful ones like in-context learning (Brown et al., 2020).
Pretrained models can represent high-level and low-level semantics, so they can represent high-level instructions and low-level actions. This semantic understanding can extend across modalities while retaining these task-learning abilities, forming combined representations of language and vision useful to robotics problems (OpenAI et al., 2023). Beyond their representation capacity, a unified interface for training and inference makes these methods particularly powerful. Multiple models across modalities can be connected and trained end-to-end through gradient descent, and pretrained representations can be improved through increases in model and data scale  (Brown et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022). Provided with a method for generating and standardizing data, a single pipeline with minimal human intervention can be used for continued improvement and deployment. One limitation of deep methods is that they require abundant data, which limits their applicability to domains where data is easy to collect. However, large-scale pretraining can significantly reduce the amount of task-specific training data required (Reed et al., 2022a; Bommasani et al., 2021).
@@ 3.1 Image and Language Subgoals
These methods use images and natural language to express subgoals. Black et al. (2023) perform pick-and-place tasks on a robot by alternating between generating images representing sub-goals using a text-guided image-editing model and executing a low-level policy conditioned on the goal images. They map language to subgoals represented as images and then map these images to end-effector positions. VLP (Du et al., 2024) transforms the input image and command into a sequence of language and image subgoals and then generates plan rollouts with tree search using a language-conditioned video generation model. It was deployed on a robot to perform tabletop arrangement tasks.
UniSim (Yang et al., 2023) trains a generative video model to predict the outcome of both high and low-level actions. Using this model, the authors train high-level planners and low-level policies in simulation and demonstrate zero-shot transfer to real-world autonomous driving tasks. Language and images are mapped to high-level skills and low-level controls, e.g., “move the gripper to (x, y)”. GAIA-1 (Hu et al., 2023) similarly leverages real-world driving data to develop a world model that generates life-like driving behavior by predicting outcomes in the video space. Coarse symbols, such as words in English, are not as rich of a representation as images. Incorporating multimodality increases the generalizability and expressiveness of the systems’ states. This allows for directly representing the visual world state without introducing an intermediate representation like natural language, scene graphs, or maps. The downside to using images is that it requires more pretraining and fine-tuning data to learn this multimodal representation.
@@ 3.2 End-Effector and Joint-State Goals
These approaches use low-level skills and parameterize actions using joint states or end-effector position, in contrast to high-level skills like “pick up” that we classify on the formal side of the spectrum. Often, they are not learned because they are sufficiently low-level, and the built-in controllers on the robot are sufficient. For example, the RT-* family of models (Brohan et al., 2023; Zitkovich et al., 2023; Padalkar et al., 2023) use an action space that specifies the arm, base, and end-of-episode token. For the arm, they use the end-effector pose of the gripper (vs. lower-level joint states or joint torques) and rely on inverse kinematics to find joint states to move to the end-effector pose.
@@
RT-1, RT-2, and RT-X (Brohan et al., 2023; Zitkovich et al., 2023; Padalkar et al., 2023) perform various language-conditioned tasks on 22 distinct robotic platforms. They collected a large and diverse dataset of demonstrations and trained a large multimodal transformer that maps language and state observations to low-level discrete end-effector controls. These methods attempt to demonstrate positive transfer, the transfer of knowledge between tasks that allows a system to perform better at both tasks simultaneously. The Open-X Embodiment dataset (Padalkar et al., 2023) release came with baselines showing this phenomenon, whereas GATO (Reed et al., 2022b) did not. Open-X Embodiment targeted end-effector positions and used models from 35M to 55B parameters, whereas GATO targeted joint states directly and used 1.2B parameters. Octo (Octo Model Team et al., 2023) also shows better performance across tasks by leveraging the Open-X Embodiment dataset while simultaneously outputting end-effector positions and joint-state control. PaLM-E (Driess et al., 2023) is an image and text multimodal model finetuned end-to-end for various multimodal reasoning tasks, including robot manipulation. This model maps visual-language input describing a scene to a limited vocabulary of actions, which are then turned into robot actions. VIMA (Jiang et al., 2023) learns a multimodal model that maps tasks specified with images and language to high-level actions parameterized by low-level end-effector controls to perform tabletop arrangement tasks on a simulated robot. One consequence of this level of abstraction is that what is learned is often specific to a robotic platform and cannot be generalized from one environment and robot to another without more data. Generating data automatically in simulation allowed it to overcome limitations on the data required to learn a meaningful representation. Future work is focused on increasing the size and diversity of the datasets to enable more generalization.
@@
Recent works in foundation models have shown tremendous improvement in generalization by leveraging large amounts of data. These methods often use symbols in the form of vector-quantized image patches to reduce the number of tokens required (Yan et al., 2023). ALOHA (Zhao et al., 2023) seeks to address the limited data for training these systems by utilizing accessible hardware for many research institutions. They then use an end-to-end foundation model that targets joint-state control of robots to execute policies on the bimanual robot platform. Li et al. (2023b) addresses domain limitations by leveraging vision and text models that are larger and trained on more data. Shridhar et al. (2023) develop a language-conditioned behavior-cloning model that encodes the inputs of a natural language command and a voxel grid to predict a collision-free pose that a motion planner can execute. Similarly, VPT (Baker et al., 2022) utilizes real demonstrations of users playing Minecraft to train better policies that target a discrete set of actions for the agent to follow. They expand their dataset by training an inverse dynamics model to label data from YouTube. Joint-state and end-effector goals are useful for simultaneously targeting multiple robotic platforms but can lack the interpretability of their output.
@@
Command Types Domains Command Examples Implementation Examples
Temporal Navigation “Move back and forth between the table and the countertop twice.” (Liang et al., 2023)
“Go to the store on Main Street, but only after visiting the bank” (Liu et al., 2023b)
“Remain on the first floor and navigate to the red room .” (Pan et al., 2023)
Manipulation “Scan for blocks and insert any found into the bin.” (Gopalan et al., 2018)
“Create a stack that contains two blocks.” (Xie et al., 2023)
Mobile Manipulation “Get Glass of Milk” (Huang et al., 2022a)
“Always take the pear and go to the tree and stay there.” (Wang et al., 2021)
“I spilled my coke. Can you bring me something to clean it up?” (Ichter et al., 2022)
“Can you bring me the drink from the table?” (Huang et al., 2022b)
“Bring the plate to the kitchen” (Hsiung et al., 2022)
“Move KeyChain1 to the box with more books” (Xie et al., 2023)
Spatiotemporal Navigation “Go to Landmark A then Landmark B, in addition, avoid Landmark C;” “Move back and forth between Object A in Room B and Object C in Room D” (Liu et al., 2023b; Pan et al., 2023)
Manipulation “Put the red block to the left of the rightmost bowl” (Liang et al., 2023)
“Pack the ring into the brown box” (Hsu et al., 2023)
“Sort fruits on the plate and bottles in the box” (Singh et al., 2023)
“Move all the blocks to different corners.” (Zeng et al., 2023a)
“May I have a cup of milk with taro?” (Li et al., 2023a)
“Stack Objects X, Y, and Z” “Push Object X to Object Y” (Black et al., 2023; Brohan et al., 2023; Driess et al., 2023; Black et al., 2023; Yang et al., 2023; Jiang et al., 2023; Liang et al., 2023; Du et al., 2024; Zitkovich et al., 2023; Padalkar et al., 2023)
Mobile Manipulation “Take the Coca-Cola can from the desk and put it in the middle of fruits on the table” (Liang et al., 2023)
“Microwave salmon” (Singh et al., 2023)
“Bring Object A to Location B;” “Open the door and proceed to the kitchen” (Ichter et al., 2022; Driess et al., 2023; Huang et al., 2022b; Hsiung et al., 2022)
“Build a house out of dirt blocks” (Wang et al., 2023)
“Put away groceries in the pantry” (Huang et al., 2022a; Li et al., 2022)
@@ 4.1 Representation
Table 1 summarizes the grounding representations used by various systems we review and how they execute their output on a robot in various simulated and real-world domains. On the formal side of the spectrum, system output ranges from more structured logic formulas and code to less structured predefined skills. On the deep side of the spectrum, methods ground language input to high-dimensional representations of joint states or end-effector poses. Some deep approaches first ground language to an intermediate representation of image or language subgoals, then low-level controls. Formal approaches to executing the system output on a robot use a planner or a code interpreter together with low-level controllers. In contrast, deep approaches use low-level controllers to produce desired robot motion. A key takeaway is using intermediate representations at different system levels, which enables the use of “off-the-self” robot modules such as SLAM and object detectors. When using these modules, there is a crisp abstraction barrier between the learning model and the robot, limiting what needs to be learned and the flexibility of the learned system. Automatically learning an extendable set of symbols and grounding those symbols (Gopalan et al., 2020) can address this limitation of formal approaches. Another key research question for formal representations is what formal language to use. Yet a formal language that is powerful and flexible enough to capture all English or other natural languages is unknown. Approaches such as RLANG (Rodriguez-Sanchez et al., 2023) expand the scope of what language can talk about (from goals and actions to observations, states, and transition functions). Integrating these formal, manually specified languages with open-class learned models is critical, for example, jointly learning skills and symbols (Konidaris et al., 2018; Konidaris, 2020). This survey only focuses on natural language, yet using multimodal input, e.g., text, audio, RGB-D images, videos, and joint trajectories, etc., to complement language (Reid et al., 2024) in solving robotic tasks is also a promising future research direction. Current state-of-the-art approaches combine multiple modalities in a single neural architecture capable of running in real-time, including speech, image, and text OpenAI (2024). This promises to enable more audio interactivity between humans and robots.
@@ 4.3 Datasets
Both approaches require datasets for learning and evaluation. Formal methods generally require parallel datasets that map natural language to structures in the formal representation or at least the ability to test if a formal representation is correct. These parallel corpora can be expensive at scale, but new LLM methods enable good performance even with much smaller datasets. A common approach is to show a trajectory of robot behavior generated from an underlying formal representation and then ask annotators to describe the trajectory in language (Gopalan et al., 2018; Patel et al., 2020). This approach can enable untrained annotators to provide parallel data. Still, it often leads to ambiguous situations because the trajectory does not overtly show the constraints present in the underlying representation. Thus, it remains an open problem to collect diverse and unambiguous language commands and their corresponding labels.
To compensate for this problem, approaches, especially less formal approaches, try to learn from large unannotated datasets or from datasets obtained from simulation. LLMs can also enable robust learning to map from natural language to formal representation using few-shot learning or fine-tuning (Liu et al., 2023b; Hsu et al., 2023; Liu et al., 2023a; Liang et al., 2023; Ichter et al., 2022). Methods that use less constrained representations, such as VIMA (Jiang et al., 2023), collect data in simulation across many scenarios and eschew a formal representation that requires annotation of any kind, instead directly train from joint trajectories, images, and video.

### v2

@@ 2.4 Predefined Skills
Recent methods have used LLMs as a planner to map language commands to a sequence of predefined skills. Skills can be learned from data or manually specified. These predefined skills are a formal representation since they are discrete and manually specified. (Even if skills are learned, they are frequently learned from human-provided demonstrations, which provide the discrete structure for the skills.) They are also action-based representations specifying actions to take rather than goals or resultant states. For example, a robot might be provided with skills such as “pick up,” “put down,” “drive to refrigerator,” “drive to microwave,” and “clean table.” Then, the challenge is to map a natural language instruction such as “Get me the apple” and “Clean the table with the sponge” to a sequence of skills. Huang et al. (2022a) iteratively prompted an LLM to decompose a high-level task specification in natural language to a sequence of action descriptions. They used sentence similarities to map the proposed action descriptions to actions available in a simulated environment where their corresponding predefined low-level controllers can be executed. SayCan Ahn et al. (2022) iteratively prompted an LLM to sequence pre-trained skills described by predefined verb phrases based on their probabilities of success from the current state to solve mobile manipulation tasks specified by natural language on a physical robot. Their approach relies on both formal representations and high-dimensional, end-to-end differentiable representations. Despite using a fixed ontology of skills, these are implemented using a deep approach. SayCan learns these skills using a multi-task value function conditioned on language. CAPE Raman et al. (2024) also used an LLM planner to sequence predefined skills. When a plan does not meet the precondition of some skill, it re-prompted the LLM with corrective feedback. In addition to using an LLM to ground language to predefined skills, Inner Monologue Huang et al. (2022b) also enabled language feedback based on pre-trained perception models for describing the scene and detecting successful execution of skills. In these methods, the specific skills are often given relatively little attention, yet having the right set of skills at the right level of granularity is critical to success. We categorize this method on the formal side of the spectrum because the skills impose significant structure. That said, it imposes less structure than goal-based methods such as logical representations, leaving the LLM to keep track of higher-level constructs such as constraints, sequencing, and conditionals.
@@ 3 Mapping from Natural Language to High-Dimensional Vectors to Actions
On the other end of the spectrum are approaches that map natural language instructions to high-dimensional, non-formal representations. These deep approaches are largely defined by data and learning, whereas formal, symbolic representations are largely human-crafted and invoke manually provided structure. The large models that enable deep and many symbolic approaches are created by self-supervised pretraining on large datasets scraped from the web. These models thereby learn a generative representation of images, text, and other modalities that capture background and task knowledge that is useful for robotic applications, for example, as in  Driess et al. (2023). By learning to model a large and varied distribution, the models also learn to perform tasks useful for robotic language grounding, including translation, semantic parsing, and broadly useful ones like in-context learning Brown et al. (2020).
Pretrained models can represent high-level and low-level semantics, so they can represent high-level instructions and low-level actions. This semantic understanding can extend across modalities while retaining these task-learning abilities, forming combined representations of language and vision useful to robotics problems. Beyond their representation capacity, a unified interface for training and inference makes these methods particularly powerful. Multiple models across modalities can be connected and trained end-to-end through gradient descent, and pretrained representations can be improved through increases in model and data scale  Brown et al. (2020); Henighan et al. (2020); Hoffmann et al. (2022); Wei et al. (2022). Provided with a method for generating and standardizing data, a single pipeline with minimal human intervention can be used for continued improvement and deployment. One limitation of deep methods is that they require abundant data, which limits their applicability to domains where data is easy to collect. However, large-scale pretraining can significantly reduce the amount of task-specific training data required Brown et al. (2020).
@@ 3.1 Image and Language Subgoals
These methods use images and natural language to express subgoals. Black et al. (2023) perform pick-and-place tasks on a robot by alternating between generating images representing sub-goals using a text-guided image-editing model and executing a low-level policy conditioned on the goal images. They map language to subgoals represented as images and then map these images to end-effector positions. VLP Du et al. (2023) transforms the input image and command into a sequence of language and image subgoals and then generates plan rollouts with tree search using a language-conditioned video generation model. It was deployed on a robot to perform tabletop arrangement tasks.
UniSim Yang et al. (2023) trains a generative video model to predict the outcome of both high and low-level actions. Using this model, the authors train high-level planners and low-level policies in simulation and demonstrate zero-shot transfer to real-world autonomous driving tasks. Language and images are mapped to high-level skills and low-level controls, e.g., “move the gripper to (x, y)”. GAIA-1 Hu et al. (2023) similarly leverages real-world driving data to develop a world model that generates life-like driving behavior by predicting outcomes in the video space. Coarse symbols, such as words in English, are not as rich of a representation as images. Incorporating multimodality increases the generalizability and expressiveness of the systems’ states. This allows for directly representing the visual world state without introducing an intermediate representation like natural language, scene graphs, or maps. The downside to using images is that it requires more pretraining and fine-tuning data to learn this multimodal representation.
@@ 3.2 End-Effector and Joint-State Goals
These approaches use low-level skills and parameterize actions using joint states or end-effector position, in contrast to high-level skills like “pick up” that we classify on the formal side of the spectrum. Often, they are not learned because they are sufficiently low-level, and the built-in controllers on the robot are sufficient. For example, the RT-* family of models Brohan et al. (2023a, b); O’Neill et al. (2024) use an action space that specifies the arm, base, and end-of-episode token. For the arm, they use the end-effector pose of the gripper (vs. lower-level joint states or joint torques) and rely on inverse kinematics to find joint states to move to the end-effector pose.
@@
RT-1, RT-2, and RT-X Brohan et al. (2023a, b); O’Neill et al. (2024) perform various language-conditioned tasks on 22 distinct robotic platforms. They collected a large and diverse dataset of demonstrations and trained a large multimodal transformer that maps language and state observations to low-level discrete end-effector controls. These methods attempt to demonstrate positive transfer, the transfer of knowledge between tasks that allows a system to perform better at both tasks simultaneously. The Open-X Embodiment dataset O’Neill et al. (2024) release came with baselines showing this phenomenon, whereas GATO Reed et al. (2022) did not. Open-X Embodiment targeted end-effector positions and used models from 35M to 55B parameters, whereas GATO targeted joint states directly and used 1.2B parameters. Octo Octo Model Team et al. (2023) also shows better performance across tasks by leveraging the Open-X Embodiment dataset while simultaneously outputting end-effector positions and joint-state control. PaLM-E Driess et al. (2023) is an image and text multimodal model finetuned end-to-end for various multimodal reasoning tasks, including robot manipulation. This model maps visual-language input describing a scene to a limited vocabulary of actions, which are then turned into robot actions. VIMA Jiang et al. (2023) learns a multimodal model that maps tasks specified with images and language to high-level actions parameterized by low-level end-effector controls to perform tabletop arrangement tasks on a simulated robot. One consequence of this level of abstraction is that what is learned is often specific to a robotic platform and cannot be generalized from one environment and robot to another without more data. Generating data automatically in simulation allowed it to overcome limitations on the data required to learn a meaningful representation. Future work is focused on increasing the size and diversity of the datasets to enable more generalization.
@@
Recent works in foundation models have shown tremendous improvement in generalization by leveraging large amounts of data. These methods often use symbols in the form of vector-quantized image patches to reduce the number of tokens required Yan et al. (2023). ALOHA Zhao et al. (2023) seeks to address the limited data for training these systems by utilizing accessible hardware for many research institutions. They then use an end-to-end foundation model that targets joint-state control of robots to execute policies on the bimanual robot platform. Li et al. (2023b) addresses domain limitations by leveraging vision and text models that are larger and trained on more data. Shridhar et al. (2023) develop a language-conditioned behavior-cloning model that encodes the inputs of a natural language command and a voxel grid to predict a collision-free pose that a motion planner can execute. Similarly, VPT Baker et al. (2022) utilizes real demonstrations of users playing Minecraft to train better policies that target a discrete set of actions for the agent to follow. They expand their dataset by training an inverse dynamics model to label data from YouTube. Joint-state and end-effector goals are useful for simultaneously targeting multiple robotic platforms but can lack the interpretability of their output.
Command Types Domains Command Examples Implementation Examples
Temporal Navigation “Move back and forth between the table and the countertop twice.” Liang et al. (2022)
“Go to the store on Main Street, but only after visiting the bank” Liu et al. (2023b)
“Remain on the first floor and navigate to the red room .” Pan et al. (2023)
Manipulation “Scan for blocks and insert any found into the bin.” Gopalan et al. (2018)
“Create a stack that contains two blocks.” Xie et al. (2023)
Mobile Manipulation “Get Glass of Milk” Huang et al. (2022a)
“Always take the pear and go to the tree and stay there.” Wang et al. (2021)
“I spilled my coke. Can you bring me something to clean it up?” Ahn et al. (2022)
“Can you bring me the drink from the table?” Huang et al. (2022b)
“Bring the plate to the kitchen” Hsiung et al. (2022)
“Move KeyChain1 to the box with more books” Xie et al. (2023)
Spatiotemporal Navigation “Go to Landmark A then Landmark B, in addition, avoid Landmark C;” “Move back and forth between Object A in Room B and Object C in Room D” Liu et al. (2023b); Pan et al. (2023)
Manipulation “Put the red block to the left of the rightmost bowl” Liang et al. (2022)
“Pack the ring into the brown box” Hsu et al. (2023)
“Sort fruits on the plate and bottles in the box” Singh et al. (2023)
“Move all the blocks to different corners.” Zeng et al. (2023a)
“May I have a cup of milk with taro?” Li et al. (2023a)
“Stack Objects X, Y, and Z” “Push Object X to Object Y” Black et al. (2023); Brohan et al. (2023a); Driess et al. (2023); Black et al. (2023); Yang et al. (2023); Jiang et al. (2023); Liang et al. (2022); Du et al. (2023); Brohan et al. (2023b); O’Neill et al. (2024)
Mobile Manipulation “Take the Coca-Cola can from the desk and put it in the middle of fruits on the table” Liang et al. (2022)
“Microwave salmon” Singh et al. (2023)
“Bring Object A to Location B;” “Open the door and proceed to the kitchen” Ahn et al. (2022); Driess et al. (2023); Huang et al. (2022b); Hsiung et al. (2022)
“Build a house out of dirt blocks” Wang et al. (2023)
“Put away groceries in the pantry” Huang et al. (2022a); Li et al. (2022)
@@ 4.1 Representation
Table 1 summarizes the grounding representations used by various systems we review and how they execute their output on a robot in various simulated and real-world domains. On the formal side of the spectrum, system output ranges from more structured logic formulas and code to less structured predefined skills. On the deep side of the spectrum, methods ground language input to high-dimensional representations of joint states or end-effector poses. Some deep approaches first ground language to an intermediate representation of image or language subgoals, then low-level controls. Formal approaches to executing the system output on a robot use a planner or a code interpreter together with low-level controllers. In contrast, deep approaches use low-level controllers to produce desired robot motion. A key takeaway is using intermediate representations at different system levels, which enables the use of “off-the-self” robot modules such as SLAM and object detectors. When using these modules, there is a crisp abstraction barrier between the learning model and the robot, limiting what needs to be learned and the flexibility of the learned system. Automatically learning an extendable set of symbols and grounding those symbols Gopalan et al. (2020) can address this limitation of formal approaches. Another key research question for formal representations is what formal language to use. Yet a formal language that is powerful and flexible enough to capture all English or other natural languages is unknown. Approaches such as RLANG Rodriguez-Sanchez et al. (2023) expand the scope of what language can talk about (from goals and actions to observations, states, and transition functions). Integrating these formal, manually specified languages with open-class learned models is critical, for example, jointly learning skills and symbols Konidaris et al. (2018); Konidaris (2020). This survey only focuses on natural language, yet using multimodal input, e.g., text, audio, RGB-D images, videos, and joint trajectories, etc., to complement language Reed et al. (2022) in solving robotic tasks is also a promising future research direction. Current state-of-the-art approaches combine multiple modalities in a single neural architecture capable of running in real-time, including speech, image, and text Reid et al. (2024). This promises to enable more audio interactivity between humans and robots.
@@ 4.3 Datasets
Both approaches require datasets for learning and evaluation. Formal methods generally require parallel datasets that map natural language to structures in the formal representation or at least the ability to test if a formal representation is correct. These parallel corpora can be expensive at scale, but new LLM methods enable good performance even with much smaller datasets. A common approach is to show a trajectory of robot behavior generated from an underlying formal representation and then ask annotators to describe the trajectory in language Gopalan et al. (2018); Patel et al. (2020). This approach can enable untrained annotators to provide parallel data. Still, it often leads to ambiguous situations because the trajectory does not overtly show the constraints present in the underlying representation. Thus, it remains an open problem to collect diverse and unambiguous language commands and their corresponding labels.
To compensate for this problem, approaches, especially less formal approaches, try to learn from large unannotated datasets or from datasets obtained from simulation. LLMs can also enable robust learning to map from natural language to formal representation using few-shot learning or fine-tuning Liu et al. (2023b); Hsu et al. (2023); Liu et al. (2023a); Liang et al. (2022); Ahn et al. (2022). Methods that use less constrained representations, such as VIMA Jiang et al. (2023), collect data in simulation across many scenarios and eschew a formal representation that requires annotation of any kind, instead directly train from joint trajectories, images, and video.